# Topology-Preserving Image Segmentation with Spatial-Aware Persistent Feature Matching

### v1

@@ Abstract
Topological correctness is critical for segmentation of tubular structures. Existing topological segmentation loss functions are primarily based on the persistent homology of the image. They match the persistent features from the segmentation with the persistent features from the ground truth and minimize the difference between them. However, these methods suffer from an ambiguous matching problem since the matching only relies on the information in the topological space. In this work, we propose an effective and efficient Spatial-Aware Topological Loss Function that further leverages the information in the original spatial domain of the image to assist the matching of persistent features. Extensive experiments on images of various types of tubular structures show that the proposed method has superior performance in improving the topological accuracy of the segmentation compared with state-of-the-art methods.
footnotetext: Our codes will be available at the time of publication.
@@ 1 Introduction
To the best of our knowledge, the only existing work that tried to tackle this issue is the Betti-Matching Loss (BMLoss) [29], which enables a more accurate persistent feature matching via an induced matching method [1]. However, due to both complexity of the algorithm and the limitations in its implementation, the computational cost of the BMLoss is very high, limiting its applicability to practical tasks-the experiments in [29] can only be run on very small datasets with compromised image resolution.
@@
In this paper, we propose a novel Spatial-Aware Topological Segmentation Loss Function (SATLoss) which leverages the spatial relationship between persistent features in the original image in addition to their global topological features (Fig. 1). While achieving comparable performance with the BMLoss, our method is significantly faster. The introduction of spatial-awareness considerably reduces the incorrect matching between persistent features. As a result, when compared with the other state-of-the-art (SOTA) methods on larger datasets, our proposed method achieves a major improvement in the topological accuracy.
@@ 3.2.1 Spatial-Aware Matching and Loss Formulation
We introduce the location of creators as a reference for the spatial location of the persistent features. To be more specific, considering a super-level filtration, we compute the normalized spatial distance between the creators (if dimension of creator is larger than 0, then use its 0-dimensional face with the largest filtration value) of each pair of persistent features in $\mathcal{D}(L)$ and $\mathcal{D}(T)$. Then, the distance is applied as a spatial weight $s_{p}$ to the distances between two points in the persistent diagram. We follow [13, 33, 4] and use the Wasserstein distance [20]:
$\vspace{-0.2cm}W_{q}(\mathcal{D}(L),\mathcal{D}(T))=[\inf_{\eta:\mathcal{D}(L)\rightarrow\mathcal{D}(T)}\sum_{p\in\mathcal{D}(L)}\left\|p-\eta(p))\right\|^{q}]^{\frac{1}{q}}$ (2)
@@
$\mathcal{L}_{topo}=\sum_{p}s^{*}_{p}([b(p)-b(\eta^{*}(p))]^{2}+[d(p)-d(\eta^{*}(p))]^{2})$ (4)
where $\eta^{*}(p)$ is the optimal matching of $p$ in $T$ by Eq. 3, $b(\cdot)$, $d(\cdot)$ are the birth and death filtration values of the persistent feature and $s^{*}_{p}$ is the corresponding spatial weight. The loss is further performed over the 0 and 1-persistent diagrams and summed. In reality, the number of persistent features in $L$ is usually larger than that in $T$ and the redundant persistent features in $L$ are matched to the diagonal (dotted line in Fig. 2(d)) of the persistent diagram. In another word:
@@
Note that we only consider the location of creators but not destroyers for two reasons: (1) There are persistent features which have infinite persistence thus have no destroyer; (2) For 0-features, the destroyer location of one feature can be close to another 0-feature, which could make the spatial-aware matching less accurate. In comparison, the creator of a 0-feature is the maximum pixel in a feature and the creator of a 1-feature appears at where the loop is closed and can be spatially more informative.
Fig. 3 shows matching examples between T and L from a converged segmentation model. From the limited examples we manually compared, it seems the Betti-matching [29] is indeed a more accurate method. But our proposed method still significantly reduces the matching error compared with the Wasserstein matching at a much smaller computational cost than the Betti-matching. Unfortunately, there is no ‘ground truth’ available to compare the matching accuracy numerically. Moreover, it is worth mentioning that although the matching on segmentation from a converged model offers a partial insight on this problem, it does not provide a comprehensive understanding of how the matching precisely influences the whole training process, which can be less straightforward to see. Accordingly, the final segmentation quality on testing images serves as the gold standard for the evaluation and empirically it is shown that a less confused matching enables better topological accuracy in the segmentation ([29], Table 1).
@@ 3.2.3 Computational Cost
Our SATLoss is more efficient than BMLoss [29] as it introduces no additional cost to computing the persistent homology of $L$ and $G$. As explained in Sec. 3.2.2, the computation of the critical cell locations is a necessary step in the optimization, whether or not we use them to perform the spatial-aware matching. Additionally, the cost of retrieving spatial location of the critical cells is also small, which simply requires mapping the critical cells back to the unsorted cubical complex, and is trivial comparing with the computation of persistent homology [7, 5]. Given an $M\times N$ image and $n=M\times N$, our SATLoss has a $\mathcal{O}(n\log(n))$ time complexity. In comparison, the BMLoss requires a lot more computationally heavy steps besides the computation of persistent homology of $L$ and $G$ and has a $\mathcal{O}(n^{3})$ time complexity. We will show in Sec. 4.5 that even when training on a very small patch size of $48\times 48$, there is a huge difference in the runtime. If extended to larger images, the difference can be even much larger.
@@ 4.1 Datasets
Table 4: Comparison of training time with BMLoss, the runtime is measured on an older computer with a Xeno E5-2650 CPU and GTX 1080Ti GPUs since the implementation of BMLoss requires an old PyTorch version which incurs a runtime error during training on newer GPU models.
@@
The images are cropped to a patch size of $48\times 48$ to further reduce the computational cost. More details on the 6 datasets can be found in the supplementary materials (SM).
@@ 4.2 Implementation Details
We adopt a consistent U-Net [26] network structure and set of hyperparameters for all experiments. An Adam optimizer is used with a weight decay of 1e-3 and an initial learning rate of 1e-3, decayed to 1e-4 at half of the total number of epochs. We train 30 epochs for Roads, 50 epochs for CREMI and CrackTree and 100 epochs for DRIVE. For all experiments, the model with the best Dice score on the validation set is used for testing. We use the best pixel loss reported in the corresponding paper for each of the baseline methods and use the BCE loss as pixel loss with our proposed method. The best weights $\lambda$ for the topological loss functions are determined by the ablation studies in Sec. 4.6. All experiments are repeated with 3 random seeds. For the smaller CrackTree and DRIVE datasets, we further adopt a 3-fold cross validation under each seeds. Our implementation is under the PyTorch framework, where the computation of persistent homology relies on the GUDHI (Geometry Understanding in Higher Dimensions) library. The GUDHI library is written in C++ and provides a Python interface.
@@ 4.5 Comparison with BMLoss
For comparison with BMLoss, we also include the clDice and TCLoss in the baseline, as the first method is the newest baseline in [29] and TCLoss is a newer method than BMLoss (which did not include BMLoss in their baseline). On another note, although the metrics are commonly computed on the image patches [13, 25, 27], in [29] they are computed over the full image. We therefore report the results on both the patch size P and full image size F. The result are shown in Table 2, Table 3 and Fig. 5. While our proposed method does not achieve the same performance level in global topological accuracy on full image size with BMLoss on the Roads-small dataset, on boarder evaluations which take the other results into consideration, our method demonstrates comparable, if not superior, performance. Furthermore, as reported in Table 4, our method achieves the results at a much smaller training time compared with BMLoss. Such a fundamental constraint of BMLoss is much less a problem of our method. It is important to note that in addition to the algorithm complexity, the coding language and implementation details can also influence the runtime. The bottom-level codes we use to compute persistent homology are written in C++, while the codes for BMLoss are written in Python. So this comparison is unfair in a strict sense. Given the complexity of the BMLoss, a reimplementation in C++ is challenging and is beyond the scope of this study. To the best of our knowledge, the Python version we used is the only and official implementation available. Therefore, despite the differences in implementation context, our proposed method still demonstrates superior computational efficiency.
@@ 6.5 Roads-small
We follow as much details as we can in [29] to prepare the Roads-small dataset. In total 134 images are selected from the original Roads dataset (mixing training, validation and testing, since we use 3-fold cross validation for this dataset). In [29], which images are selected is not provided so we pick image by ourselves. It is difficult to describe the selected images unless providing a long list here but the general rule is to select images with more complexed road networks and less blank regions. The selected images are then divided randomly into training, validation and testing for each fold. The images are then downsampled from $1500\times 1500$ to $375\times 375$. Again, details of downsampling are not provided in [29], so we use (the same for C.Elegan-small) bicubic interpolation for all images and labels and then binarized the downsampled labels by thresholding them with a value of 0.5. No further manual correction is made to the downsampled labels. Before cropping, the images are padded to the right and bottom to $384\times 384$. Then, each image is cropped to 64 $48\times 48$ patches. In total, there are 6400 for training, 640 for validation and 1536 for testing.
@@ 9 Additional Results
The ablation results on the other datasets are provided in Fig. 8, Fig. 9, Fig. 10, Fig. 11, Fig. 12, Fig. 13. Due to limited resources and for consistency between datasets, the ablation experiments were only run on the first fold. In addition, we provide more qualitative results in Fig. 14 for our main experiments. In Fig. 7, we also show more qualitative comparison with on the Roads-small dataset where our method fails to outperform the BMLoss in some metrics. In Fig. 15, we show more persistent feature matching results.

### v2

@@ Abstract
Topological correctness is critical for segmentation of tubular structures. Existing topological segmentation loss functions are primarily based on the persistent homology of the image. They match the persistent features from the segmentation with the persistent features from the ground truth and minimize the difference between them. However, these methods suffer from an ambiguous matching problem since the matching only relies on the information in the topological space. In this work, we propose an effective and efficient Spatial-Aware Topological Loss Function that further leverages the information in the original spatial domain of the image to assist the matching of persistent features. Extensive experiments on images of various types of tubular structures show that the proposed method has superior performance in improving the topological accuracy of the segmentation compared with state-of-the-art methods [^1] Our codes will be available at the time of publication..
@@ 1 Introduction
To the best of our knowledge, the only existing work that tried to tackle this issue is the Betti-Matching Loss (BMLoss) [30], which enables a more accurate persistent feature matching via an induced matching method [1]. However, it requires computing the refined barcodes [30] and imposes a $\mathcal{O}(n^{3})$ time complexity, making this method excessively slow, limiting its applicability to practical tasks. As a matter of fact, the experiments in [30] can only be run on very small toy datasets with compromised image resolution.
@@
In this paper, we propose a novel Spatial-Aware Topological Segmentation Loss Function (SATLoss) which takes a different approach and directly leverages the spatial relationship between persistent features in the original image in addition to their global topological features (Fig. 1) for persistent feature matching. While achieving comparable performance with the BMLoss, our method is significantly faster. The introduction of spatial-awareness considerably reduces the incorrect matching between persistent features. As a result, when compared with the other state-of-the-art (SOTA) methods on larger datasets, our proposed method achieves a major improvement in the topological accuracy.
@@ 3.2.1 Spatial-Aware Matching and Loss Formulation
We introduce the location of creators as a reference for the spatial location of the persistent features. To be more specific, considering a super-level filtration, we compute the normalized (x, y coordinates divided by image height and width, respectively) spatial distance between the creators (if dimension of creator is larger than 0, then use its 0-dimensional face with the largest filtration value) of each pair of persistent features in $\mathcal{D}(L)$ and $\mathcal{D}(T)$. Then, the distance is applied as a spatial weight $s_{p}$ to the distances between two points in the persistent diagram. We follow [13, 34, 4] and use the Wasserstein distance [20]:
$W_{q}(\mathcal{D}(L),\mathcal{D}(T))=[\inf_{\eta:\mathcal{D}(L)\rightarrow\mathcal{D}(T)}\sum_{p\in\mathcal{D}(L)}\left\|p-\eta(p))\right\|^{q}]^{\frac{1}{q}}$ (2)
@@
$\mathcal{L}_{topo}=\sum_{p\in\mathcal{D}(L)}s^{*}_{p}([b(p)-b(\eta^{*}(p))]^{2}+[d(p)-d(\eta^{*}(p))]^{2})$ (4)
where $\eta^{*}(p)$ is the optimal matching of $p$ in $T$ by Eq. 3, $b(\cdot)$, $d(\cdot)$ are the birth and death filtration values of the persistent feature and $s^{*}_{p}=\left\|c_{b}(p)-c_{b}(\eta^{*}(p))\right\|^{2}$ is the corresponding spatial weight. Eq. 4 is a reformulation of Eq. 3 given the matching (spatially-weighted optimal transport) results and the distances in $\mathcal{D}$ reformulated by the birth and death values of the topological feature. The loss is further performed over the 0 and 1-persistent diagrams and summed. In reality, the number of persistent features in $L$ is usually larger than that in $T$ and the redundant persistent features in $L$ are matched to the diagonal (Fig. 1; dotted line in Fig. 2(d)) of the persistent diagram. In another word:
@@
Note that we only consider the location of creators but not destroyers for two reasons: (1) There are persistent features which have infinite persistence thus have no destroyer; (2) For 0-features, the destroyer location of one feature can be close to another 0-feature, which could make the spatial-aware matching less accurate. In comparison, the creator of a 0-feature is the maximum pixel in a feature and the creator of a 1-feature appears at where the loop is closed and can be spatially more informative. Additionally, in implementation we apply a lower bound of 0.05 to $s_{p}$ to prevent the spatial weight from being over-dominating in the matching.
Fig. 3 shows matching examples between T and L from a converged segmentation model (more detailed examples in supplementary material (SM)). From the limited examples we manually compared, it seems the Betti-matching [30] is indeed a more accurate method. But our proposed method still significantly reduces the matching error compared with the Wasserstein matching at a much smaller computational cost than the Betti-matching. Unfortunately, there is no ‘ground truth’ available to compare the matching accuracy numerically. Moreover, it is worth mentioning that although the matching on segmentation from a converged model offers a partial insight on this problem, it does not provide a comprehensive understanding of how the matching precisely influences the whole training process, which can be less straightforward to see. Accordingly, the final segmentation quality on testing images serves as the gold standard for the evaluation and empirically it is shown that a less confused matching enables better topological accuracy in the segmentation ([30], Table 1).
@@ 3.2.3 Computational Cost
Given an $M\times N$ image and $n=M\times N$, our SATLoss has a $\mathcal{O}(n\log(n))$ time complexity, bottle-necked by the step of sorting the cubical complex in computing persistent homology (PH). The cost of retrieving spatial location of the critical cells is small, which simply requires mapping the critical cells back to the unsorted cubical complex, and is trivial comparing with the computation of PH [7, 5]. In comparison, to address the same PH mismatch problem, the BMLoss requires computing the refined barcodes [30], where the x-axis values of the barcodes are the indices in the cubical complex instead of filtration values and the computation also requires a reversing of the 1-cubes, which result in a $\mathcal{O}(n^{3})$ time complexity. We will show in Sec. 4.5 that even when using a very small patch size of $48\times 48$, there is a huge difference in the runtime. If extended to larger images, the difference can be even much larger (see SM).
@@ 4.1 Datasets
Table 4: Comparison of training time with BMLoss, the runtime is measured on an older computer with a Xeno E5-2650 CPU and GTX 1080Ti GPUs since the implementation of BMLoss requires an old PyTorch version which incurs a runtime error during training on newer GPU models. Full comparison is shown in SM.
@@
The images are cropped to a patch size of $48\times 48$ to further reduce the computational cost. More details on the 6 datasets can be found in SM.
@@ 4.2 Implementation Details
Since the topic is the loss function, we follow [28, 13, 30, 12, 14] and adopt a consistent, general U-Net [27] network structure and set of hyperparameters for all experiments. An Adam optimizer is used with a weight decay of 1e-3 and an initial learning rate of 1e-3, decayed to 1e-4 at half of the total number of epochs. We train 30 epochs for Roads, 50 epochs for CREMI and CrackTree and 100 epochs for DRIVE. We use the best pixel loss reported in the corresponding paper for each of the baseline methods and use the BCE loss as pixel loss with our proposed method. The best weights $\lambda$ for the topological loss functions are determined by the ablation studies in Sec. 4.6. All experiments are repeated with 3 random seeds. For the smaller CrackTree and DRIVE datasets, we further adopt a 3-fold cross validation under each seed. In each independent random experiment, the model with the best Dice score on the validation set is used for testing and the average test scores and their stds over these experiments are reported. Our implementation is under the PyTorch framework, where the computation of persistent homology relies on the GUDHI (Geometry Understanding in Higher Dimensions) library [25].
@@ 4.5 Comparison with BMLoss
For comparison with BMLoss, we also include the clDice and TCLoss in the baseline, as the first method is the newest baseline in [30] and TCLoss is a newer method than BMLoss (which did not include BMLoss in their baseline). On another note, although the metrics are commonly computed on the image patches [13, 26, 28], in [30] they are computed over the full image. We therefore report the results on both the patch size P and full image size F. The result are shown in Table 2, Table 3 and Fig. 5. While our proposed method does not achieve the same performance level in global topological accuracy on full image size with BMLoss on the Roads-small dataset, on boarder evaluations which take the other results into consideration, our method demonstrates comparable, if not superior, performance. Furthermore, as reported in Table 4, our method achieves the results at a much smaller training time compared with BMLoss. Such a fundamental constraint of BMLoss is much less a problem of our method. It is important to note that in addition to the algorithm complexity, the coding language and implementation details can also influence the training time. The bottom-level codes in the GUDHI library we use to compute persistent homology are written in C++, while the codes for BMLoss are all written in Python. So this comparison is unfair in a strict sense. Given the complexity of the BMLoss, a reimplementation in C++ is challenging and is beyond the scope of this study. To the best of our knowledge, the Python version we used is the only and official implementation available. Therefore, despite the differences in implementation context, our proposed method still demonstrates superior computational efficiency.
@@ 6.5 Roads-small
We follow as many details as we can in [30] to prepare the Roads-small dataset. In total 134 images are selected from the original Roads dataset (mixing training, validation and testing, since we use 3-fold cross validation for this dataset). In [30], which images are selected is not provided so we pick image by ourselves. It is difficult to describe the selected images unless providing a long list here but the general rule is to select images with more complexed road networks and less blank regions. The selected images are then divided randomly into training, validation and testing for each fold. The images are then downsampled from $1500\times 1500$ to $375\times 375$. Again, details of downsampling are not provided in [30], so we use (the same for C.Elegan-small) bicubic interpolation for all images and labels and then binarized the downsampled labels by thresholding them with a value of 0.5. No further manual correction is made to the downsampled labels. Before cropping, the images are padded to the right and bottom to $384\times 384$. Then, each image is cropped to 64 $48\times 48$ patches. In total, there are 6400 for training, 640 for validation and 1536 for testing.
@@ 9 Additional Results
Table 5 shows the full training time comparison. The ablation results on the other datasets are provided in Fig. 8, Fig. 9, Fig. 10, Fig. 11, Fig. 12, Fig. 13. Due to limited resources and for consistency between datasets, the ablation experiments were only run on the first fold. In addition, we provide more qualitative results in Fig. 14 for our main experiments. In Fig. 7, we also show more qualitative comparison with on the Roads-small dataset where our method fails to outperform the BMLoss in some metrics. In Fig. 15, we show more persistent feature matching results. In Fig. 16, we provide an illustrative example of the motivation of the spatial-aware matching, how it works during the matching process and how it helps with the topological accuracy of the segmentation.
Table 5: Training time comparison (i9-13900K+RTX 4090)
Hours #iter clDice [26] He. Ours BMLoss(Est.)
Roads 52k 2.9 44 43 44 $>$20000
CREMI 26k 1.0 16 16 16 $>$8000
CrackTree 3.7k 0.23 3.5 3.5 3.5 $>$1400
DRIVE 2.4k 0.03 0.22 0.23 0.23 $>$100
@@
Figure 16: An illustrative example of how the matching works and how spatial-aware matching helps improve topological accuracy in the segmentation. The lower three tables show the spatial weight between each pair of persistent features, the cost matrix used by vanilla Wasserstein matching in previous methods, and the cost matrix after weighted by the spatial weight used by our proposed SATLoss, respectively. The green boxes refer to correct matchings and the red boxes refer to incorrect matchings. D refers to the diagonal of the persistent diagram. Note that by Wasserstein distance (optimal transport) the matching plan giving the least overall cost is selected, to each feature there is no guarantee that the minimal-cost matching is selected. The upper figures shows the matching results and the gradient behavior (see Sec. 3.2.2 for explanation). At the destroyer of the feature g (yellow cross), by vanilla Wasserstein matching the pixel value is pushed down (since matched with GT features), making feature a and g more separated (which should be connected). Similarly, at the destroyer of the feature b (pink cross), the gradient pushes the value up (since matched with diagonal), making feature b and c connected (which should be separated). In comparison, using the proposed spatial-aware matching, at the destroyer of feature g the pixel value is pushed up, making it more connected to feature a; at the destroyer of feature b the pixel value is pushed down, making it more separated from feature c. Therefore, allowing the segmentation model to make topologically more correct predictions. Similar process occurs repeatedly at different locations on the image in the optimization iterations.